# Information Discovery in e-Commerce

### v1

@@ 5.2.1 Relevance-based metrics
Relevance-aware evaluation metrics are in various information retrieval domains [schutze2008introduction], including in e-commerce search. It is common to see studies compute evaluation metrics based on the top 100 items retrieved by each e-commerce search model [ai2017learning, van2016learning, van2018mix]. Mean average precision (MAP), hit ratio (HR), mean reciprocal rank (MRR), and normalized discounted cumulative gain (NDCG) are four widely-used relevance-aware metrics [van2018mix, wu2018turning]. All these metrics are also widely applied in various information retrieval domains [schutze2008introduction]. Average precision (AP) computes the average value of Precision over the interval from 00 to $1$. Given $k$ candidate items, $AP@k=\sum_{k=1}^{n}{P@k\cdot rel(k)}$, where $P@k$ refers to Precision@k; $rel(k)$ indicates $1$ if the $k$-th item is relevant and 00 otherwise $rel(k)=0$. Based on that, MAP calculates the mean of AP for all queries:
@@ 5.2.2 Revenue-aware metrics
GMV indicates the total income amount transacted from merchandise sales, whereas the overall revenue generated for the e-commerce site is proportional to the GMV [wu2018turning]. Revenue-aware metrics are applied to e-commerce search evaluation to evaluate how the methods can improve the actual revenue of search sessions. To calculate the average revenue for every impression in each e-commerce search session, wu2018turning proposed the average revenue metric, $Avg.Rev$ $(i,q)$, for a query-item pair $(i,q)$ as follows:
@@
wu2018turning proposed a metric named $Rev@k$ to evaluate the revenue in e-commerce rankers. $Rev@k$ calculates the average revenues that a prediction algorithm would generate for each session. Specifically, $Rev@k$ is calculated as follows:
@@
where $\rho$ is the ranking order and $r_{s}\leq k$ denotes the top-$k$ ranked positions in the session $s$. $r_{s}^{-1}$ denotes the corresponding item at the position $r_{s}$, $price(i)$ indicates the price of item $i$, while $\Phi$ denotes an event of purchase. Based on $Rev@k$, it is able to evaluate the revenue influence of the candidate rankers.
Empirical studies have been performed in benchmark e-commerce search datasets to find differences between relevance-based metrics and revenue-aware metrics [wu2018turning]. Relevance-based metrics primarily measure the success of retrieving relevant data from user logs, while revenue-aware metrics provide a clearer understanding of how ranking methods influence actual revenue in e-commerce scenarios. This difference shows that, although relevance is important, incorporating revenue-aware metrics offers more practical insights into the financial impact of search ranking methods within e-commerce settings. The lack of annotated real-world benchmarks poses a challenge for conducting large-scale empirical studies in e-commerce scenarios. To address this issue, recent studies have started using both synthetic and semi-synthetic datasets in their experiments, allowing for more controlled analysis while approximating real-world conditions [xu2022product].
@@ 5.2.3 User engagement metrics
To evaluate the quality of short-term user-system interactions, short-term engagement metrics about instant clicks, purchases, and dwell time are used in e-commerce search evaluation. As we discussed in Section 4.1 and Section 4.1.3, the click-through rate (CTR) and conversion rate (CVR) are the two most widely applied metrics to evaluate instant click and purchase prediction. Meanwhile, dwell time and bounce rate are the two main metrics used in short-term post-click evaluations [lalmas2015promoting].
As we discussed in Section 3.2.2, long-term user engagement reflects the user’s desire to stay on the e-commerce portal longer and use the service repeatedly [zou2018drlunderreview], i.e., the “stickiness.” Long-term user engagements measure versatile user behaviors based on a very large number of environmental interactions. Multiple long-term engagement metrics have been applied in previous studies. wu2017returning employ cumulative clicks over time to estimate the long-term interactions between the user and the system. zou2018drlunderreview applied $3$ evaluation metrics to evaluate the long-term user interaction behaviors: (1) average clicks per session: the average cumulative number of clicks over a user visit; (2) average depth per session: the average browsing depth that the users interact with the recommender agent; and (3) average return time: the average revisiting days between. In multilingual scenarios, a transformed query converts a secondary language query into a semantically equivalent query in the primary language, allowing it to fully utilize the existing search engines’ abilities. To evaluate transformed queries in multilingual e-commerce search, a set of behavior metrics based on user engagement specific to the existing search system was developed. Using these metrics, a query transformation system was built and tested both offline and through online A/B tests on the Amazon platform, showing improvements in the multilingual search experience for customers [hu2020query].
@@ 5.3.1 Vocabulary gap
Users’ shopping lists often differ from the product information maintained by retailers [nurmi2008product]. duan2013probabilistic found that while query languages such as SQL can be successfully applied to search in these product databases, their usage is difficult for non-experienced end users. In direct search scenarios in e-commerce platforms, most consumers formulate queries using characteristics of the product they are interested in (e.g., terms that describe the product’s categories, brands, and shops, etc.). Hence, it is common to see a mismatch in e-commerce search between queries and product representations, where different tokens are used to describe the same product, i.e., the vocabulary gap problem [van2016learning]. To address this problem, early studies focus on rewriting verbose queries in e-commerce search [bendersky2009analysis, xue2010improving, singh2012rewriting]. Selecting a subset of the original query (i.e., “sub-query”) has been shown to be effective for improving these queries. xue2010improving proposed to formally model the distribution of sub-queries, where the sub-query selection procedure is modeled as a sequential labeling problem. A conditional random field model was applied to track the local and global dependencies between query words. singh2012rewriting presented techniques to reduce long queries to effective shorter ones that lack superfluous terms. Specifically, the authors proposed a system that provides high quality product recommendations for null queries, where time-based relevance feedback is used to improve the fidelity of rewrites. However, these approaches focus only on the query space and overlook critical information from the product space and the connection between the two spaces [van2016learning].
@@ 5.3.2 Representation-based matching
In early work, approaches to the task of entity finding have been applied to address the vocabulary gap problem between queries and products, where products are viewed as retrievable entities [balog2010entity, gade2016overview, van2016learning]. However, there are two problems in entity finding for e-commerce search [de2007overview]: First, entity finding retrieves entities of a particular type from multi-domain knowledge bases, whereas e-commerce search systems operate within a single but dynamic domain. Second, queries in e-commerce search contain a lot of free-form text [rowley2000product], whereas in entity finding most queries are semi-structured with relational constraints [balog2010entity].
@@
To address the above two problems of matching optimization, representation learning methods have been applied to obtain better representations for each text associated with products by encoding both the query and the product title into single embedding vectors. Representation learning helps by flexibly adapting to the dynamic nature of e-commerce domains. It also effectively encodes the free-form text of queries and product descriptions into a common semantic space. Many numerous neural text matching methods have been developed [onal-neural-2018, mitra2017introduction, lin-2021-pretrained]. DSSM is one of the earliest deep learning-based models in text matching, in which each text is vectorized separately by a five-layer network [huang2013learning]; CDSSM replaces the full connection layer with a convolution layer and a pooling layer to generate text vectors [shen2014learning]. hu2014convolutional proposed ARC-I, where convolution operations represent two concatenated texts for matching using a linear transformation. CNTN also adopts convolution neural network to represent two texts, and it proposes the neural tensor network to model the similarities between two texts [qiu2015convolutional]. MVLSTM obtains representations for each text and adopts an interactive method to measure similarities between two texts using 3 similarity operations, i.e., cosine, bilinear, and tensor layer [wan2016deep]. As there are important differences between web search and e-commerce search, learning query and product representations is not a solved problem. Hence, to discriminate products based on textual descriptions, the importance of learning semantic representations of products was soon realized [demartini2009vector, van2016unsupervised].
@@
van2016learning proposed an unsupervised distributed representation learning approach, namely latent semantic entities (LSE), to learn a unidirectional mapping between words and entities, as well as distributed representations of both words and entities. Given a set of entities, the authors assume that each entity has a set of associated documents. LSE then learns a function that maps a sequence of words in the query from the vocabulary to an entity vector space. Thereafter, cosine similarity is applied to calculate a relevance score between candidate entities and the query. In Fig. 5.1, we see how entities are then ranked according to the projected query.
@@
In addition to learning the semantic representation of products, e-commerce search, due to its specific task nature, also has multiple methods for optimizing matching results. Substitutable products are those that are interchangeable in e-commerce platforms [Wang2018Improving]. The substitutability relation among products can be determined in multiple ways. van2018mix found that product substitutability relations can facilitate the retrieval of relevant products impacted by the vocabulary gap problem, i.e., product substitutability can be integrated into product search either extrinsically or intrinsically. The authors proposed a two stage framework to combine textual matching method (i.e., the LSE model) and substitutability to infer representations of queries and products. Unlike previous work on entity representation learning, the authors integrated relations among entities within the latent semantic space inference.
@@
Users often browse multiple search results pages and make comparisons before purchase. Relevance feedback (RF) approaches were proposed to extract the relevant topic [jin2013interactive]. However, the mismatch problem between queries and products still existed in the task of multi-page e-commerce search. bi2019study analyzed different context dependency assumptions in multiple search result pages, and proposed a context-aware embedding model to capture different types of dependency. The authors introduced three different types of context dependencies: (i) long-term context dependency, (ii) short-term context dependency, and (iii) long-short-term context dependency. Given these three types of context dependencies, a context-aware embedding model was proposed. By assuming that the users’ preferences are associated with their implicit feedback, the embedding model, namely CEM, captures user preferences from their clicked items, which are implicit positive signals.
@@
Another important aspect of e-commerce search is query reformulations by users. Based on query logs of eBay’s search engine, hirsch2020query proposed the first large-scale and in-depth study of users’ query reformulations in e-commerce search. The authors analyzed many aspects of search sessions composed of query reformulations, e.g., the number of reformulations and the distribution of their types, changes of search results pages as a result of the reformulations, clicks and purchases. An approach was proposed to predict if a query will be reformulated in an e-commerce search session. The authors find that post-retrieval features and query performance predictors contribute the most to the prediction of reformulation. By incorporating these features, the accuracy in predicting whether users will reformulate their queries can be significantly enhanced. Based on the attention mechanism, the MMAN model was proposed to enhance query representation by extending category information and includes three main modules: self-matching, char-level matching, and semantic-level matching [Yuan_2023]. Experiments have shown that these modules improve query representation, effectively handle long-tail queries, and achieve better semantic disambiguation.
@@
In recent years, graph neural networks (GNNs) [scarselli2008graph, kipf2016semi] have been applied to e-commerce search to help infer query and product embeddings. GNNs derive node representations by aggregating features appearing in the neighborhood. niu2020dual proposed a GNN-based method to learn representations of users and shops in e-commerce search. Specifically, the authors proposed a dual hierarchical graph attention network for e-commerce search. A heterogeneous graph was constructed to perform graph-based representation learning for both shops and queries, which includes both first-order and second-order proximities from various user interactions in e-commerce. The authors verified that their proposed method can help to relieve the semantic gap between user queries and shop names by borrowing item neighbor title text. Furthermore, the proposed neighbor proximity loss provides strong additional guidance for learning graph topological structure. The results of large-scale offline evaluations and online A/B tests demonstrate the significant superiority of this approach. chang2021extreme transferred the matching problem as an extreme multi-label classification problem [yu2022pecos], aiming to tag input instances (i.e., queries) with the most relevant output labels of products. The authors proposed a tree-based sparse linear model with n-gram TF-IDF features to augment the diversity of the matching results. For multi-lingual search scenarios in e-commerce, lu2021graph proposed a graph-based model with a graph convolution layer to fill the vocabulary gap.
@@
Transformer-based pre-trained models like BERT [kenton2019bert] utilize stacked encoder layers that rely on a self-attention mechanism. This mechanism allows each word in a sequence to attend to all other words, facilitating capturing richer semantic relations throughout the sequence. chang2020pre conducted a comprehensive study applying a pre-trained dual-tower matching model across multiple retrieval tasks, demonstrating the effectiveness of such models in enhancing retrieval performance. BERT and BERT-like pre-trained models has been applied in product search [peeters2020intermediate, liu2022towards, qiu2022pre, wang2023learning]. qiu2022pre applied dual-tower pre-training strategies to optimize both user intent detection and embedding retrieval in e-commerce search. liu2022towards examined the performance of multiple pre-training embedding methods and observed that query representation learning remains a bottleneck compared to product representation learning when using these semantic search training objectives. Recently, large-scale pre-trained language models, such as GPT-3 [brown2020language], also demonstrate promising performance across several benchmarks [kim2022ask].
@@ 5.3.3 Interaction-based matching
However, encoding queries and products in representation-based matching methods are independent of each other. These individual queries and products into the fixed-dimensional vectors may lose the fine-grained matching information. To tackle this challenge, interaction-based matching was proposed. This kind of method first matches different parts of the query with different parts of the document and then aggregates the partial evidence of relevance. In contrast to representation-based matching methods, interaction-based approaches usually build an interaction matrix between two documents and optimize it for matching [hu2014convolutional, pang2016text]. This interaction-based matching approach was applied in web-based retrieval at first. hu2014convolutional built an interaction matrix to conduct several convolution and pooling operations to extract matching features. guo2016deep mentioned three factors in relevance matching: exact matching signals, query term importance, and diverse matching requirements, and designed the architecture of their deep matching model. Similarly, the model of MatchPyramid constructs an interaction matrix to capture matching patterns [pang2016text]. mitra2017learning employed both distributed representations and local representations to obtain the final matching score. Subsequently, researchers proposed a series of matching approaches specifically for e-commerce built upon existing interaction-based approaches. guo2019matchzoo proposed a model based on the MatchZoo. It is meant for short-text matching and replaces the matching histogram with a top-k max pooling layer. li2020deep proposed a product matching model, namely PMM, to make use of the information contained in titles and attributes of products, respectively. PMM consists of two modules, a product title matching module, and a product attributes matching module. The former computes similarities between product titles, whereas the latter measures similarities between product attribute set. bi2019leverage proposed an end-to-end context-aware embedding model that can incorporate both long-term and short-term contexts to predict purchased items, unlike most approaches that focus on relevance feedback.

### v3

@@ 5.2.1 Relevance-based metrics
Relevance-aware evaluation metrics are in various information retrieval domains [schutze2008introduction], including in e-commerce search. It is common to see studies compute evaluation metrics based on the top 100 items retrieved by each e-commerce search model [ai2017learning, van2016learning, van2018mix]. Mean average precision (MAP), hit ratio (HR), mean reciprocal rank (MRR), and normalized discounted cumulative gain (NDCG) are four widely-used relevance-aware metrics [van2018mix, wu2018turning]. These metrics are also widely applied in various information retrieval domains [schutze2008introduction]. Average precision (AP) computes the average value of Precision over the interval from 00 to $1$. Given $k$ candidate items, $AP@k=\sum_{k=1}^{n}{P@k\cdot rel(k)}$, where $P@k$ refers to Precision@k; $rel(k)$ indicates $1$ if the $k$-th item is relevant and 00 otherwise $rel(k)=0$. Based on that, MAP calculates the mean of AP for all queries:
@@ 5.2.2 Revenue-aware metrics
GMV indicates the total income amount transacted from merchandise sales, whereas the overall revenue generated for the e-commerce site is proportional to the GMV [wu2018turning]. Revenue-aware metrics are applied to e-commerce search evaluation to evaluate how the methods can improve the actual revenue of search sessions. To calculate the average revenue for every impression in each e-commerce search session, wu2018turning introduce the average revenue metric, $Avg.Rev$ $(i,q)$, for a query-item pair $(i,q)$ as follows:
@@
wu2018turning specify a metric named $Rev@k$ to evaluate the revenue in e-commerce rankers. $Rev@k$ calculates the average revenues that a prediction algorithm would generate for each session. Specifically, $Rev@k$ is calculated as follows:
@@
where $\rho$ is the ranking order and $r_{s}\leq k$ denotes the top-$k$ ranked positions in the session $s$. $r_{s}^{-1}$ denotes the corresponding item at the position $r_{s}$, $price(i)$ indicates the price of item $i$, while $\Phi$ denotes a purchase event. Based on $Rev@k$, it is able to evaluate the revenue influence of the candidate rankers.
Empirical studies have been performed in benchmark e-commerce search datasets to find differences between relevance-based metrics and revenue-aware metrics [wu2018turning]. Relevance-based metrics primarily measure the success of retrieving relevant data from user logs, while revenue-aware metrics provide a clearer understanding of how ranking methods influence actual revenue in e-commerce scenarios. This difference shows that, although relevance is important, incorporating revenue-aware metrics offers more practical insights into the financial impact of search ranking methods within e-commerce settings.
The lack of annotated real-world benchmarks poses a challenge for conducting large-scale empirical studies in e-commerce scenarios. To address this issue, recent studies have started using both synthetic and semi-synthetic datasets in their experiments, allowing for more controlled analysis while approximating real-world conditions [xu2022product].
@@ 5.2.3 User engagement metrics
To evaluate the quality of short-term user-system interactions, short-term engagement metrics about instant clicks, purchases, and dwell time are used in e-commerce search evaluation. As we discussed in Section 4.1 and 4.1.3, the click-through rate (CTR) and conversion rate (CVR) are the two most widely applied metrics to evaluate instant click and purchase prediction. Meanwhile, dwell time and bounce rate are the two main metrics used in short-term post-click evaluations [lalmas2015promoting].
As we discussed in Section 3.2.2, long-term user engagement reflects the user’s desire to stay on the e-commerce portal longer and use the service repeatedly [zou2018drlunderreview], i.e., the “stickiness.” Long-term user engagements measure versatile user behaviors based on a very large number of environmental interactions. Multiple long-term engagement metrics have been applied in previous studies. wu2017returning employ cumulative clicks over time to estimate the long-term interactions between the user and the system. zou2018drlunderreview apply three evaluation metrics to evaluate the long-term user interaction behaviors: (i) average clicks per session: the average cumulative number of clicks over a user visit; (ii) average depth per session: the average browsing depth that the users interact with the recommender agent; and (iii) average return time: the average revisiting days between. In multilingual scenarios, a transformed query converts a secondary language query into a semantically equivalent query in the primary language, allowing it to fully use the search engines’ abilities. To evaluate transformed queries in multilingual e-commerce search, a set of behavior metrics based on user engagement specific to the existing search system was developed. Using these metrics, a query transformation system has been built and tested both offline and through online A/B tests on the Amazon platform, showing improvements in the multilingual search experience for customers [hu2020query].
@@ 5.3.1 Vocabulary gap
Users’ shopping lists often differ from the product information maintained by retailers [nurmi2008product]. duan2013probabilistic find that while query languages such as SQL can be successfully applied to search in these product databases, their usage is difficult for non-experienced end users. In direct search scenarios in e-commerce platforms, most consumers formulate queries using characteristics of the product they are interested in (e.g., terms that describe the product’s categories, brands, and shops, etc.). Hence, it is common to see a mismatch in e-commerce search between queries and product representations, where different tokens are used to describe the same product, i.e., the vocabulary gap problem [van2016learning]. To address this problem, early studies focus on rewriting verbose queries in e-commerce search [bendersky2009analysis, xue2010improving, singh2012rewriting]. Selecting a subset of the original query (i.e., “sub-query”) has been shown to be effective for improving these queries. xue2010improving formally model the distribution of sub-queries, where the sub-query selection procedure is modeled as a sequential labeling problem. A conditional random field model was applied to track the local and global dependencies between query words. singh2012rewriting present techniques to reduce long queries to effective shorter ones that lack superfluous terms. The authors describe a system that provides high quality product recommendations for null queries, where time-based relevance feedback is used to improve the fidelity of rewrites. However, these approaches focus only on the query space and overlook critical information from the product space and the connection between the two spaces [van2016learning].
@@ 5.3.2 Representation-based matching
In early work, approaches to the task of entity finding have been applied to address the vocabulary gap problem between queries and products, where products are viewed as retrievable entities [balog2010entity, gade2016overview, van2016learning]. There are two problems in entity finding for e-commerce search [de2007overview]. First, entity finding retrieves entities of a particular type from multi-domain knowledge bases, whereas e-commerce search systems operate within a single but dynamic domain. Second, queries in e-commerce search contain a lot of free-form text [rowley2000product], whereas in entity finding most queries are semi-structured with relational constraints [balog2010entity].
@@
To address the above two problems of matching optimization, representation learning methods have been applied to obtain better representations for each text associated with products by encoding both the query and the product title into single embedding vectors. Representation learning helps by flexibly adapting to the dynamic nature of e-commerce domains. It also effectively encodes the free-form text of queries and product descriptions into a common semantic space. Numerous neural text matching methods have been developed [onal-neural-2018, mitra2017introduction, lin-2021-pretrained]. DSSM is one of the earliest deep learning-based models in text matching, in which each text is vectorized separately by a five-layer network [huang2013learning]; CDSSM replaces the full connection layer with a convolution layer and a pooling layer to generate text vectors [shen2014learning]. hu2014convolutional proposed ARC-I, where convolution operations represent two concatenated texts for matching using a linear transformation. CNTN also adopts convolution neural network to represent two texts, and it proposes the neural tensor network to model the similarities between two texts [qiu2015convolutional]. MVLSTM obtains representations for each text and adopts an interactive method to measure similarities between two texts using 3 similarity operations, i.e., cosine, bilinear, and tensor layer [wan2016deep]. As there are important differences between web search and e-commerce search, learning query and product representations is not a solved problem. Hence, to discriminate products based on textual descriptions, the importance of learning semantic representations of products was soon realized [demartini2009vector, van2016unsupervised].
@@
van2016learning propose an unsupervised distributed representation learning approach, namely latent semantic entities (LSE), to learn a unidirectional mapping between words and entities, as well as distributed representations of both words and entities. Given a set of entities, the authors assume that each entity has a set of associated documents. LSE then learns a function that maps a sequence of words in the query from the vocabulary to an entity vector space. Thereafter, cosine similarity is applied to calculate a relevance score between candidate entities and the query. In Fig. 5.1, we see how entities are then ranked according to the projected query.
@@
In addition to learning semantic representations of products, e-commerce search, due to its specific task nature, also has multiple methods for optimizing matching results. Substitutable products are those that are interchangeable in e-commerce platforms [Wang2018Improving]. The substitutability relation among products can be determined in multiple ways. van2018mix find that product substitutability relations can facilitate the retrieval of relevant products impacted by the vocabulary gap problem, i.e., product substitutability can be integrated into product search either extrinsically or intrinsically. The authors propose a two-stage framework to combine a textual matching method (the LSE model) with substitutability to infer representations of queries and products. Unlike previous work on entity representation learning, the authors integrate relations among entities within the latent semantic space inference.
@@
Users often browse multiple search results pages and make comparisons before purchase. Relevance feedback (RF) approaches have been proposed to extract the relevant topic [jin2013interactive]. However, the mismatch problem between queries and products still exists in the task of multi-page e-commerce search. bi2019study analyze different context dependency assumptions in multiple search result pages, and propose a context-aware embedding model to capture different types of dependency. The authors introduced three types of context dependency: (i) long-term context dependency, (ii) short-term context dependency, and (iii) long-short-term context dependency. Given these three types of context dependency, a context-aware embedding model is proposed. By assuming that the users’ preferences are associated with their implicit feedback, the embedding model, namely CEM, captures user preferences from their clicked items, which are implicit positive signals.
@@
Another important aspect of e-commerce search concerns query reformulations by users. Based on query logs of eBay’s search engine, hirsch2020query offer a large-scale and in-depth study of users’ query reformulations in e-commerce search. The authors analyze many aspects of search sessions composed of query reformulations, e.g., the number of reformulations and the distribution of their types, changes of search results pages as a result of the reformulations, clicks and purchases. An approach is proposed to predict if a query will be reformulated in an e-commerce search session. The authors find that post-retrieval features and query performance predictors contribute the most to the prediction of reformulation. By incorporating these features, the accuracy in predicting whether users will reformulate their queries can be significantly enhanced. Based on an attention mechanism, the MMAN model is proposed to enhance query representations by extending category information; it includes three main modules: self-matching, char-level matching, and semantic-level matching [Yuan_2023]. Experiments show that these modules improve query representation, effectively handle long-tail queries, and achieve better semantic disambiguation.
@@
Graph neural networks (GNNs) [scarselli2009graph, kipf2016semi] have been applied to e-commerce search to help infer query and product embeddings. GNNs derive node representations by aggregating features appearing in the neighborhood. niu2020dual put forward a GNN-based method to learn representations of users and shops in e-commerce search. The authors propose a dual hierarchical graph attention network for e-commerce search. A heterogeneous graph is constructed to perform graph-based representation learning for both shops and queries, which includes both first-order and second-order proximities from various user interactions in e-commerce. The proposed method can help to relieve the semantic gap between user queries and shop names by borrowing item neighbor title text. The proposed neighbor proximity loss provides strong additional guidance for learning graph topological structure. A large-scale offline evaluation and online A/B tests demonstrate the significant superiority of this approach. chang2021extreme transfer the matching problem into an extreme multi-label classification problem [yu2022pecos], aiming to tag input instances (i.e., queries) with the most relevant output labels of products. The authors suggest a tree-based sparse linear model with n-gram TF-IDF features to augment the diversity of the matching results. For multi-lingual search scenarios in e-commerce, lu2021graph detail a graph-based model with a graph convolution layer to fill the vocabulary gap.
@@
Transformer-based pre-trained models like BERT [kenton2019bert] use stacked encoder layers that rely on a self-attention mechanism. BERT and BERT-like pre-trained models have been applied in product search [peeters2020intermediate, liu2022towards, qiu2022pre, wang2023learning]. qiu2022pre apply dual-tower pre-training strategies to optimize both user intent detection and embedding retrieval in e-commerce search. liu2022towards examine the performance of multiple pre-training embedding methods and observed that query representation learning remains a bottleneck compared to product representation learning when using these semantic search training objectives. Large-scale pre-trained language models, such as GPT-3 [brown2020language], also demonstrate promising performance across several benchmarks [kim2022ask].
@@ 5.3.3 Interaction-based matching
Encoding queries and products in representation-based matching methods are independent of each other. Mapping individual queries and products into fixed-dimensional vectors may lose fine-grained matching information. To tackle this challenge, interaction-based matching has been proposed. This kind of method first matches different parts of the query with different parts of the document and then aggregates the partial evidence of relevance. In contrast to representation-based matching methods, interaction-based approaches usually build an interaction matrix between two documents and optimize it for matching [hu2014convolutional, pang2016text]. Interaction-based matching has first been applied in web-based retrieval. hu2014convolutional build an interaction matrix to conduct several convolution and pooling operations to extract matching features. guo2016deep mention three factors in relevance matching – exact matching signals, query term importance, and diverse matching requirements – and design the architecture of their deep matching model. Similarly, MatchPyramid constructs an interaction matrix to capture matching patterns [pang2016text]. mitra2017learning employ both distributed representations and local representations to obtain the final matching score. Follow-up research has proposed a series of matching approaches specifically for e-commerce. guo2019matchzoo introduce a model based on MatchZoo. It is meant for short-text matching and replaces the matching histogram with a top-$k$ max pooling layer. li2020deep describe a product matching model, PMM, to make use of the information contained in titles and attributes of products. PMM consists of a product title matching module, and a product attributes matching module. bi2019leverage detail an end-to-end context-aware embedding model that can incorporate both long-term and short-term contexts to predict purchased items, unlike most approaches that focus on relevance feedback.